OpenStudy (crashonce):

fThe marked price of a calculator is x. 25% of this price is profit. If the calculator is sold at discount of 10%, what is the profit?

OpenStudy (crashonce):

@perl

OpenStudy (danjs):

The full price is x. 25 % of that is profit or 0.25x = Profit. Then it is marked down by 10 % to 0.25(x-0.1x) = new profit

OpenStudy (danjs):

25 percent of 90 percent of the full price

OpenStudy (perl):

ok the marked price is the price before the discount

OpenStudy (crashonce):

so its 0.225

OpenStudy (crashonce):

is it? cos it isnt in the multiple choice

OpenStudy (crashonce):

? @perl

OpenStudy (crashonce):

?

OpenStudy (crashonce):

@mathmath333 can u get the answer

OpenStudy (mathmath333):

wait

OpenStudy (mathmath333):

by assumption method let the marked price be \(100\) hence profit is \(25\) hence cost price is \(100-25=75\) discounted price is \(90\) \(new~~ profit %=\dfrac{90-75}{75}\times 100\)

OpenStudy (mathmath333):

\(new~~ profit \%= \dfrac{90-75}{75}\times 100\)

OpenStudy (crashonce):

the answers are 0.21x, 0.12x, 0.18x, 0.15x, 0.1x

OpenStudy (mathmath333):

here marked price =\(100=x\)

OpenStudy (mathmath333):

as far as i know the discount is given on marked price so it should be \(20\%\) ie \(0.2x\)

OpenStudy (perl):

if we use mathmath's method 100 dollars was the original marked price of the 100 dollars, 75 dollars is cost, 25 dollars is profit. what happens when we reduce the price by 10% the cost doesnt change since you have to pay your distributor, so we lose profit

OpenStudy (mathmath333):

wait \(profit =selling~price-cost~price\\ =90-75\\ =15\)

OpenStudy (mathmath333):

@ perl it is asking about profit not profit %\)

OpenStudy (mathmath333):

so it should be \(0.15x\)

OpenStudy (perl):

right

OpenStudy (perl):

because x = 100 was the original marked price

OpenStudy (perl):

in our example

OpenStudy (crashonce):

?

OpenStudy (perl):

here is another way to do it, algebraically Using the equation profit = selling price - cost price Let x = marked price before 10% discount of marked price: profit = x - .75x after 10% discount of marked price: profit = .9x - .75x

OpenStudy (perl):

cost price is the price the seller has to pay to the producer or the distributor, to get the stuff to sell. That cost is fixed , the seller can't change that

OpenStudy (crashonce):

ye ok thanks

OpenStudy (mathmath333):

marked price -discount=selling price

OpenStudy (crashonce):

ye i get it thanks both